# Pretrained Reversible Generation as Unsupervised Visual Representation Learning

### v2

@@ 4.3.1 Better Pre-training Lead to Better Fine-tuning?
where $\mathcal{H}(X_{0})$ is the (constant) entropy of the data distribution and $\mathcal{H}(X_{1})$ is the (constant) entropy of the Gaussian distribution. Hence, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ faithfully reflects the mutual information $\mathcal{I}(X_{0},X_{1})$ during pre-training. In contrast, during fine-tuning, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ no longer represents $\mathcal{I}(X_{0},X_{1})$, because $\mathcal{H}(X_{1})$ is no longer the entropy of a Gaussian distribution. Nonetheless, the quantity $\mathcal{I}(X_{0},X_{1})+\mathcal{H}(X_{1})\;-\;\mathcal{H}(X_{0})$ still captures the change in $\mathcal{I}(X_{0},X_{1})$ to a large extent. We compute log-probability density $\log p(x_{0}|x_{1})$ using the adjoint method or Neural ODEs [11] for computational feasibility:
@@
For efficient computation of $\log p(x_{0}|x_{1})$, we employ the Hutchinson trace estimator [28, 22]:
@@ 4.3.2 What is a Reasonable Fine-tuning Strategy?
Figure 5: Reverse generation from $x_{0}$ to $x_{1}$ with full pre-training and fine-tuning. See Appendix D for more details.
@@
Additionally, on CIFAR-10, initiating fine-tuning from later stages of the sampling trajectory ($x_{1/4}$ to $x_{1}$) results in progressively better performance. On TinyImageNet, optimal results are achieved when fine-tuning starts from $x_{1/2}$ to $x_{1}$. This difference likely stems from the complexity of the datasets: more complex datasets require stronger feature extraction, necessitating longer trajectory lengths for effective fine-tuning. Further experiments exploring optimal fine-tuning strategies are detailed in the ablation studies.
@@ 4.3.3 Is the Model a Continuous Feature Extractor?
Residual networks, such as ResNet-50 [23], achieve feature extraction by constructing discrete-time methods through a series of composite transformations: $\mathbf{h}_{t+1}=\mathbf{h}_{t}+f(\mathbf{h}_{t},\theta_{t})$. In contrast, we construct the feature extractor using the ODE specified by the neural network, namely $\frac{d\mathbf{h}(t)}{dt}=f(\mathbf{h}(t),t,\theta)$. Tabs. 1 and 2 demonstrate that although we trained using a discrete sample length of $t_{\text{span}}=20$, we can evaluate the model using any reasonable inference steps. Furthermore, feature extraction is not limited to a specific training point (such as the midpoint); extracting features within its vicinity (within 20%) still achieves high performance on downstream tasks. These findings validate our method can serve as a continuous feature extractor.
@@
WideResNet-$28$-$10$ [70] $36$ $96.3$
ResNeXt-$29$-$16*64$d [66] $68$ $96.4$
@@
GLOW [19] N/A $84.0$
Energy model [18] N/A $92.9$
SBGC [73] N/A $95.0$
HybViT [67] $43$ $96.0$
DDAE [65] $36$ $97.2$
@@
HybViT [67] $43$ $56.7$
DDAE [65] $40$ $69.4$
@@
Method Param. (M) Acc. (%)
Discriminative methods
ViT-L/$16$ ($384^{2}$) [17] $307$ $76.5$
ResNet-$152$ ($224^{2}$) [23] $60$ $77.8$
Swin-B ($224^{2}$) [39] $88$ $83.5$
Generative methods
HybViT ($32^{2}$) [67] $43$ $53.5$
DMSZC-DiTXL2 ($256^{2}$) [34] $338$ $77.5$
iGPT-L ($48^{2}$) [10] $1362$ $72.6$
[…]
PRG-GVP-onlyPretrain ($64^{2}$) $122$ $20.18$
PRG-GVP-XL ($64^{2}$) $122$ $77.84$
PRG-ICFM-XL ($64^{2}$) $122$ $78.12$
PRG-OTCFM-XL ($64^{2}$) $122$ $78.13$
Table 5: Comparison on the ImageNet dataset with various algorithms. All results are reported from their original papers, the values in parentheses indicate the input image size.
4.4 Main Results
4.4.1 Performance on Image Classification
To comprehensively evaluate our method, we conducted experiments on three image classification datasets, as detailed in Secs. 4.3.3, 4.3.3 and 4.3.3. Our methods achieved accuracies of 97.59%, 71.12%, and 78.1% on CIFAR-10, Tiny-ImageNet, and ImageNet, respectively, surpassing all existing generative approaches, as well as supervised methods like WideResNet-28[70]. However, these results still trail behind recent transformer-based supervised learning architectures like SwinTransformer [39]. We hypothesize that integrating transformer architectures and leveraging VAEs to handle high input resolutions could further enhance the performance. This avenue will be explored in future work.
4.4.2 Out-of-Distribution Robustness
To address the degradation in out-of-distribution (OOD) due to common image corruptions or adversarial perturbations [73], data augmentations and adversarial training are typically employed. However, recent studies [73, 8, 34] have indicated that generative-based classifiers, without requiring additional data, often exhibit superior OOD robustness. As illustrated in Sec. 4.4.2 and Sec. C.3, our method showcases remarkable robustness on these two datasets with only simple data augmentation.
Model CIFAR-10 CIFAR-10-C
Clean Corr Severity All w/ Noise
Adversarial Training
PGD *[43] $93.91$ $83.08$ $\downarrow 10.83$ $82.10$ $\downarrow 11.81$
PLAT * [64] $94.73$ $88.28$ $\downarrow 6.45$ 88.5688.5688.5688.566.17absent6.17$\downarrow$ 6.17↓ 6.17
Noise Injection
RSE *[37] $95.59$ $77.86$ $\downarrow 17.73$ N/A
ENResNet *[64] $83.33$ $74.34$ $\downarrow 8.99$ N/A
Data Augment
AugMix § [25] $95.83$ $89.09$ $\downarrow 6.74$ $88.71$ $\downarrow 7.12$
AutoAug *[15] $95.61$ $85.37$ $\downarrow 10.24$ 76.4776.4776.4776.4719.14absent19.14$\downarrow$ 19.14↓ 19.14
Generative Methods
SBGC [73] $95.04$ $76.24$ $\downarrow 18.80$ 75.3875.3875.3875.3819.66absent19.66$\downarrow$ 19.66↓ 19.66
PDE+ [68] $95.59$ $89.11$ $\downarrow 6.48$ 85.5985.5985.5985.5910.00absent10.00$\downarrow$ 10.00↓ 10.00
[…]
PRG-GVP-S $97.35$ $91.21$ $\downarrow 6.14$ $92.17$ $\downarrow 5.18$
PRG-ICFM-S $97.59$ $92.13$ $\downarrow 5.46$ $93.07$ $\downarrow 4.52$
PRG-OTCFM-S $97.65$ $92.26$ $\downarrow 5.39$ $93.10$ $\downarrow 4.55$
Table 6: (OOD: Image Corruptions) Comparison on CIFAR-10-C, including Noise, Blur, Weather, and Digital corruptions. Results are referenced from original papers or [68, 73]. * denotes ResNet-18 as the base model, while § indicates ResNeXt-29.
Method Param. (M) Cifar Acc. Tiny Acc.
Classic self-supervised learning methods
MAE (ViT-B/16) [40] $86$ $96.5$ $76.5$
SimCLR Res-50-4x [12] $375$ $98.6$ N/A
Generative methods
DDAE-DiT-XL2 [65] $338$ $98.4$ $77.8$
iGPT-L [10] $1362$ $99.0$ N/A
[…]
PRG-SiT-XL2 $338$ $98.72$ $78.33$
Table 7: Comparison of models pretrained on ImageNet-1k and transferred (fine-tuning) to CIFAR-10 and Tiny-ImageNet.
4.4.3 Transferring Features
A main goal of the pre-training/fine-tuning paradigm is to learn transferrable features. We hypothesize that advancements like powerful pretrained models in the generative model community can contribute to our method. To validate this, we conducted experiments by transferring the pretrained SiT-XL [42] model to two other datasets. Since SiT provides only class-conditional checkpoints, we adopted an unconditional approach by setting the label parameter to null. Specifically, we fine-tuned the model for 28 epochs on the CIFAR-10 dataset and for 45 epochs on the Tiny-ImageNet dataset. As demonstrated in Sec. 4.4.2, our method shows superior transfer learning performance. The results suggest that our algorithm benefits from larger datasets and the latent space of generative models.
4.5 Ablation Studies
In this section, we conduct three ablation studies to further evaluate the applicability of our method. Additional experiments, including analyses of loss type, ODE solver choice, dual-task supervision, and $t_{\text{span}}$, are provided in Appendix C.
4.5.1 Comparison with Generation without pre-training
@@
4.5.2 Effect of $\beta$ on Accuracy
During pre-training, we introduced $\beta$ in Eq. 3 to enhance the mutual information lower bound while balancing the downstream task objective. Table 9 shows PRG scores across datasets for different $\beta$ values.
$\beta$ 1 10 100
Acc. (CIFAR/Tiny) $0.973$/$0.712$ $0.974$/$0.711$ $0.965$/$0.673$
Table 9: ($\beta$ values) Accuracy scores of PRG with varying $\beta$ values on the CIFAR-10 and Tiny-ImageNet datasets.
4.5.3 Generative Model Type
We explored various generative model types, particularly focusing on path selection. To assess the impact of these paths on our method, we conducted experiments with three model types on the CIFAR-10 and Tiny-ImageNet datasets. Following [38], we quantified the straightness of various continuously differentiable trajectories. Interestingly, while all three model variants achieved comparable performance (Tab. 10), models with higher straightness required fewer pre-training steps to attain equivalent performance levels and demonstrated potential for superior outcomes.
Generative Model Type GVP ICFM OTCFM
CIFAR-10 (Accuracy) $97.35$ $97.59$ $97.65$
First to $97$% (Epoch) $162$ $135$ $128$
Straightness $7.36$ $0.34$ $0.11$
Tiny-ImageNet (Accuracy) $70.98$ $71.12$ $71.33$
First to $70$% (Epoch) $157$ $125$ $106$
Straightness $6.25$ $0.54$ $0.15$
Table 10: (Generative Model Type) Performance of different generative model types on the CIFAR-10 and Tiny-ImageNet datasets.
4.5.4 Scaling Up Network Parameters
We investigated the impact of scaling up U-Net models of varying network parameter sizes [48]. Table 11 provides a comparative analysis following fine-tuning on the CIFAR-10 and Tiny-ImageNet datasets. The findings suggest that increasing the model size results in a modest enhancement of final classification accuracy.
Model Param (M) CIFAR Acc. Tiny Acc.
PRG-ICFM-B $32$ $97.35$ $70.94$
PRG-ICFM-S $42$ $97.59$ $71.12$
PRG-ICFM-L $66$ $97.75$ $71.30$
PRG-ICFM-XL $122$ $97.88$ $71.80$
Table 11: (Scaling Capability) Performance gains of increasing U-Net model sizes on the CIFAR-10 and Tiny-ImageNet datasets.
5 Conclusion and Limitation
In this paper, we introduce a novel perspective on using pretrained reversible generative models as unsupervised visual representation learners. We systematically investigate the necessary designs of the two-stage pre-training and fine-tuning paradigm. Theoretical analysis supports the efficacy of the first-stage generative pre-training, while empirical verifications provide insights for designing the fine-tuning pipeline. Leveraging these simple yet effect techniques and findings, our method achieves state-of-the-art performance in image classification tasks using generative models. Additional experiments, such as those on OOD problems and transfer learning scenarios, further validate the robustness and other advantages of our approach. However, there remains room for improvement, including the integration of large foundation pretrained generative models from the open-source community (e.g., [52]) and optimizing the training epoch/speed for faster adaptation to downstream tasks.
Appendix A Why PRG is effective?
According to the manifold assumption [46, Chapter 20], data lies on a low-dimensional manifold $\mathcal{M}$ of intrinsic dimension $d^{*}$, which is significantly smaller than the ambient dimension $D$. Although, in theory, intermediate latents encode the same information as the original data, using these latents as inputs for downstream tasks is more meaningful. A well-trained flow model effectively extracts data from the low-dimensional manifold $\mathcal{M}$ and lifts it to the ambient space $\mathbb{R}^{D}$, ensuring that every sample in $\mathbb{R}^{D}$ remains semantically meaningful. In contrast, directly sampling from the original data space does not necessarily preserve semantic coherence. For instance, generating a $64\times 64$ image by sampling from a $64\times 64$ multivariate Gaussian distribution would typically result in a meaningless noise-like image resembling a snowflake pattern.
Moreover, since the flow trajectories of an ordinary differential equation (ODE) do not intersect [11], points within a given region remain confined, thereby preserving topological relationships throughout the transformation process.
The effectiveness of fine-tuning depends significantly on whether the model undergoes pre-training beforehand. If a classifier is trained directly using the ODE architecture without pre-training, the model must learn meaningful intermediate latents solely through a supervised loss. This approach often fails, as the model struggles to discover good latent representations in the absence of prior knowledge. Our experiments (Sec. 4.5.1) also confirm that this method performs poorly.
@@ A.2 Model-agnosticity
Our method follows the model-agnostic principle as defined in Appendix D.5 of [56], where different architectures achieve comparable encoding quality through flow matching. This indicates that architectural constraints stem from implementation choices rather than the framework itself.
Notably, while early diffusion models primarily relied on U-Nets for practical stability, recent work [3] has demonstrated that flow matching can be successfully achieved with minimal architectural requirements, such as shallow skip connections. In this work, we achieve state-of-the-art performance across both U-Net (Sec. 4.4.1) and Transformer (Sec. 4.4.3) architectures, further validating the generality of our approach.
@@ B.1 Training Process Details
Tabs. 1 and 2 list the hyperparameters used for both pre-training and fine-tuning across the CIFAR-10, Tiny-ImageNet, and ImageNet datasets.
@@ C.2 ODE Solver Type
During fine-tuning, we evaluated different ODE solvers for the reverse process: Euler (first-order), Midpoint (second-order via midpoint evaluations), RK4 (fourth-order Runge-Kutta), and Dopri5 (adaptive step sizes with a fifth-order method). Tabs. 6 and 7 compares their performance on the Cifar-10, Tiny-ImageNet dataset. The results show no significant performance differences, underscoring the method’s consistent effectiveness across various solvers.
@@ C.3 Details of Out-of-Distribution Experiments
PGD [43] $51.08$ $33.46$ $\downarrow 17.62$ $24.00$ $\downarrow 27.08$
PLAT [31] $51.29$ $37.92$ $\downarrow 13.37$ $29.05$ $\downarrow 22.24$
@@
RSE [37] $53.74$ $27.99$ $\downarrow 25.75$ $18.92$ $\downarrow 34.82$
ENResNet [64] $49.26$ $25.83$ $\downarrow 23.43$ $19.01$ $\downarrow 30.25$
@@
AugMix [25] $52.82$ $37.74$ $\downarrow 15.08$ $28.66$ $\downarrow 24.16$
AutoAug [15] $52.63$ $35.14$ $\downarrow 17.49$ $25.36$ $\downarrow 27.27$
@@ C.5 Transfer Experiment on Dual-Task Datasets
Following the approach described in Sec. 4.4.3, we transfer the powerful text-to-image model SiT-XL [42] from the community. However, instead of directly applying it to downstream classification tasks, we leverage its diffusion trajectory for feature extraction across multiple tasks. Specifically, we set the diffusion trajectory to 30 steps, extracting features from the 10-th reverse diffusion step for CIFAR classification and the 20-th step for Tiny-ImageNet.

### v3

@@ 4.3.1 Better Pre-training Lead to Better Fine-tuning?
where $\mathcal{H}(X_{0})$ is the (constant) entropy of the data distribution and $\mathcal{H}(X_{1})$ is the (constant) entropy of the Gaussian distribution. Hence, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ faithfully reflects the mutual information $\mathcal{I}(X_{0},X_{1})$ during pre-training. In contrast, during fine-tuning, $\mathbb{E}_{p(x_{1},x_{0})}[\log p(x_{0}\mid x_{1})]$ no longer represents $\mathcal{I}(X_{0},X_{1})$, because $\mathcal{H}(X_{1})$ is no longer the entropy of a Gaussian distribution. Nonetheless, the quantity $\mathcal{I}(X_{0},X_{1})+\mathcal{H}(X_{1})\;-\;\mathcal{H}(X_{0})$ still captures the change in $\mathcal{I}(X_{0},X_{1})$ to a large extent. We compute log-probability density $\log p(x_{0}|x_{1})$ using the adjoint method or Neural ODEs [Chen2018] for computational feasibility:
@@
For efficient computation of $\log p(x_{0}|x_{1})$, we employ the Hutchinson trace estimator [Hutchinson1989, Grathwohl2018]:
@@ 4.3.2 What is a Reasonable Fine-tuning Strategy?
Figure 5: Reverse generation from $x_{0}$ to $x_{1}$ with full pre-training and fine-tuning.
@@
Additionally, on CIFAR-10, initiating fine-tuning from later stages of the sampling trajectory ($x_{1/4}$ to $x_{1}$) results in progressively better performance. On TinyImageNet, optimal results are achieved when fine-tuning starts from $x_{1/2}$ to $x_{1}$. This difference likely stems from the complexity of the datasets: more complex datasets require stronger feature extraction, necessitating longer trajectory lengths for effective fine-tuning. For an in-depth analysis, see Sec. A.1.1; complementary experiments on optimal fine-tuning strategies are provided in Appendix C.
@@ 4.3.3 Is the Model a Continuous Feature Extractor?
Residual networks, such as ResNet-50 [he2016deep], achieve feature extraction by constructing discrete-time methods through a series of composite transformations: $\mathbf{h}_{t+1}=\mathbf{h}_{t}+f(\mathbf{h}_{t},\theta_{t})$. In contrast, we construct the feature extractor using the ODE specified by the neural network, namely $\frac{d\mathbf{h}(t)}{dt}=f(\mathbf{h}(t),t,\theta)$. Tabs. 1 and 2 demonstrate that although we trained using a discrete sample length of $t_{\text{span}}=20$, we can evaluate the model using any reasonable inference steps. Furthermore, feature extraction is not limited to a specific training point (such as the midpoint); extracting features within its vicinity (within 20%) still achieves high performance on downstream tasks. These findings validate our method can serve as a continuous feature extractor.
@@
WideResNet-$28$-$10$ [zagoruyko2016wide] $36$ $96.3$
ResNeXt-$29$-$16*64$d [xie2017aggregated] $68$ $96.4$
@@
GLOW [fetaya2019understanding] N/A $84.0$
Energy model [duvenaud2020your] N/A $92.9$
SBGC [zimmermann2021score] N/A $95.0$
HybViT [yang2022your] $43$ $96.0$
DDAE [xiang2023denoising] $36$ $97.2$
@@
HybViT [yang2022your] $43$ $56.7$
DDAE [xiang2023denoising] $40$ $69.4$
@@
Supplementary Material
Appendix A Why PRG is effective?
What a model can generate, it can understand: pixel-perfect reconstruction means the visual breadcrumbs are all there. It just needs a bit of fine-tuning to tie the generative pathway to the labels. Raw-image baselines often fall prey to ”shortcut learning”. Because neighboring feature pixels have heavily overlapping receptive fields, the network may rely on trivial cues and end up with weak representations. In contrast, PRG reuses the pretrained reverse generative pathway, recycling each ODE-solver output as the input for the next iteration. This loop drives feature extraction at a semantic level without losing low-level detail and the injected noise perturbations make the representations significantly more robust. Sec. 4.3.2 show that every point along the trajectory possesses strong representational capacity, and the OOD experiments also provide compelling evidence for this claim.
According to the manifold assumption [murphy2023probabilistic, Chapter 20], data lies on a low-dimensional manifold $\mathcal{M}$ of intrinsic dimension $d^{*}$, which is significantly smaller than the ambient dimension $D$. Although, in theory, intermediate latents encode the same information as the original data, using these latents as inputs for downstream tasks is more meaningful. A well-trained flow model effectively extracts data from the low-dimensional manifold $\mathcal{M}$ and lifts it to the ambient space $\mathbb{R}^{D}$, ensuring that every sample in $\mathbb{R}^{D}$ remains semantically meaningful. In contrast, directly sampling from the original data space does not necessarily preserve semantic coherence. For instance, generating a $64\times 64$ image by sampling from a $64\times 64$ multivariate Gaussian distribution would typically result in a meaningless noise-like image resembling a snowflake pattern.
Moreover, since the flow trajectories of an ordinary differential equation (ODE) do not intersect [Chen2018], points within a given region remain confined, thereby preserving topological relationships throughout the transformation process.
The effectiveness of fine-tuning depends significantly on whether the model undergoes pre-training beforehand. If a classifier is trained directly using the ODE architecture without pre-training, the model must learn meaningful intermediate latents solely through a supervised loss. This approach often fails, as the model struggles to discover good latent representations in the absence of prior knowledge. Our experiments (Sec. A.1.1) also confirm that this method performs poorly.
A.1.1 Comparison with Generation without pre-training
@@ A.2 Model-agnosticity
Our method follows the model-agnostic principle as defined in Appendix D.5 of [song2020score], where different architectures achieve comparable encoding quality through flow matching. This indicates that architectural constraints stem from implementation choices rather than the framework itself.
Notably, while early diffusion models primarily relied on U-Nets for practical stability, recent work [bao2023all] has demonstrated that flow matching can be successfully achieved with minimal architectural requirements, such as shallow skip connections. In this work, we achieve state-of-the-art performance across both U-Net (LABEL:sec:unet) and Transformer (LABEL:sec:trans) architectures, further validating the generality of our approach.
@@ B.1 Training Process Details
Tabs. 2 and 3 list the hyperparameters used for both pre-training and fine-tuning across the CIFAR-10, Tiny-ImageNet, and ImageNet datasets.
@@ C.2 ODE Solver Type
During fine-tuning, we evaluated different ODE solvers for the reverse process: Euler (first-order), Midpoint (second-order via midpoint evaluations), RK4 (fourth-order Runge-Kutta), and Dopri5 (adaptive step sizes with a fifth-order method). Tabs. 7 and 8 compares their performance on the Cifar-10, Tiny-ImageNet dataset. The results show no significant performance differences, underscoring the method’s consistent effectiveness across various solvers.
@@ C.3 Details of Out-of-Distribution Experiments
PGD [madry2017towards] $51.08$ $33.46$ $\downarrow 17.62$ $24.00$ $\downarrow 27.08$
PLAT [kireev2022effectiveness] $51.29$ $37.92$ $\downarrow 13.37$ $29.05$ $\downarrow 22.24$
@@
RSE [liu2018towards] $53.74$ $27.99$ $\downarrow 25.75$ $18.92$ $\downarrow 34.82$
ENResNet [wang2020enresnet] $49.26$ $25.83$ $\downarrow 23.43$ $19.01$ $\downarrow 30.25$
@@
AugMix [hendrycks2019augmix] $52.82$ $37.74$ $\downarrow 15.08$ $28.66$ $\downarrow 24.16$
AutoAug [cubuk2019autoaugment] $52.63$ $35.14$ $\downarrow 17.49$ $25.36$ $\downarrow 27.27$
@@ C.5 Transfer Experiment on Dual-Task Datasets
Following the approach described in LABEL:sec:trans, we transfer the powerful text-to-image model SiT-XL [ma2024sit] from the community. However, instead of directly applying it to downstream classification tasks, we leverage its diffusion trajectory for feature extraction across multiple tasks. Specifically, we set the diffusion trajectory to 30 steps, extracting features from the 10-th reverse diffusion step for CIFAR classification and the 20-th step for Tiny-ImageNet.